Your investment club has only two stocks in its portfolio. $40,000 is invested in a stock with a beta of 0.7, and $80,000 is invested in a stock with a beta of 1.8. What is the portfolio's beta? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Total investment=(40,000+80,000)=$120,000

Portfolio beta=Respective beta*Respective investment weight

=(40,000/120,000*0.7)+(80,000/120,000*1.8)

which is equal to

=1.43(Approx).